# An upper bound for the number of smooth values of a polynomial and its applications

### v1

@@ 3. An inductive process
There exists a constant $C(\alpha)>0$ depending only on $\alpha$ such that
$\operatorname{N}(\mathfrak{b}_{n})\leq C(\alpha)$
holds for any $n\in\mathbb{Z}$.
Proof.
Take a prime ideal $\mathfrak{p}$ of $K$ and $v\geq 1$ such that $\mathfrak{p}^{v}\mid\mathfrak{b}_{n}$. We consider an upper bound for the norm of $\mathfrak{p}^{v}$. Let $L$ denote the Galois closure of $K$ over $\mathbb{Q}$. For a rational prime $p$ with $\mathfrak{p}\mid p$, we write the factorization of $p$ in $L$ as
$p\mathcal{O}_{L}=\left(\mathfrak{P}_{1}\cdots\mathfrak{P}_{g}\right)^{e},$
where $\mathfrak{P}_{1},\ldots,\mathfrak{P}_{g}$ are prime ideals of $L$ with $N(\mathfrak{P}_{k})=p^{f}$. Note that at least one of $e\geq 2$ and $f\geq 2$ is true. Indeed, if $e=f=1$, then $p$ splits completely in $L$. This implies that $p$ splits completely also in $K$, but $\mathfrak{p}\in P_{K}$ follows in this case. Next, we factorize the prime ideal $\mathfrak{p}$ as
(5.1) $\mathfrak{p}\mathcal{O}_{L}=\left(\mathfrak{P}_{k_{1}}\cdots\mathfrak{P}_{k_{m}}\right)^{\eta},$
where $k_{1},\ldots,k_{m}\in\{1,\ldots,g\}$ and $1\leq\eta\leq e$. Denote by $M_{j}$ the decomposition field of $\mathfrak{P}_{k_{j}}$ over $\mathbb{Q}$, and define $\mathfrak{q}_{j}=\mathfrak{P}_{k_{j}}\cap\mathcal{O}_{M_{j}}$. Then $\mathfrak{P}_{k_{j}}$ is the unique prime ideal of $L$ lying over $\mathfrak{q}_{j}$. More precisely, we see that
(5.2) $\mathfrak{q}_{j}\mathcal{O}_{L}=\mathfrak{P}_{k_{j}}^{e}\quad\text{and}\quad\operatorname{N}(\mathfrak{q}_{j})=p$
by the standard properties of decomposition fields. Put $v\eta=eq+r$ with $q\geq 0$ and $0\leq r<e$. From (5.1) and (5.2), we obtain
(5.3) $\mathfrak{p}^{v}\mathcal{O}_{L}$ $=\left(\mathfrak{P}_{k_{1}}\cdots\mathfrak{P}_{k_{m}}\right)^{v\eta}$
$=\mathfrak{q}_{1}^{q}\mathcal{O}_{L}\cdots\mathfrak{q}_{m}^{q}\mathcal{O}_{L}\cdot\mathfrak{P}_{k_{1}}^{r}\cdots\mathfrak{P}_{k_{m}}^{r}.$
Hence, taking the norms, we have
(5.4) $\operatorname{N}(\mathfrak{p}^{v})^{[L:K]}=\operatorname{N}(\mathfrak{q}_{1}^{q})^{d}\cdots\operatorname{N}(\mathfrak{q}_{m}^{q})^{d}\cdot\operatorname{N}(\mathfrak{P}_{k_{1}})^{r}\cdots\operatorname{N}(\mathfrak{P}_{k_{m}})^{r},$
where $d=[L:M_{j}]=ef$. Denote by $\Delta_{L/\mathbb{Q}}$ the absolute value of the discriminant of $L$ over $\mathbb{Q}$. If $e\geq 2$, then $p\mid\Delta_{L/\mathbb{Q}}$ is satisfied. In this case, $\operatorname{N}(\mathfrak{P}_{k_{j}})^{r}=p^{fr}\leq\Delta_{L/\mathbb{Q}}^{ef}$ holds. If $e=1$, then $\operatorname{N}(\mathfrak{P}_{k_{j}})^{r}=1$ since $r$ must be equal to 00. As a result, we obtain
(5.5) $\operatorname{N}(\mathfrak{P}_{k_{1}})^{r}\cdots\operatorname{N}(\mathfrak{P}_{k_{m}})^{r}\leq\Delta_{L/\mathbb{Q}}^{efm}\leq\Delta_{L/\mathbb{Q}}^{[L:\mathbb{Q}]}$
due to $[L:\mathbb{Q}]=efg$. Next, we choose $\beta\in\mathcal{O}_{K}$ and $c\in\mathbb{Z}_{\geq 1}$ so that $\alpha=\beta/c$. Note that we can determine them uniquely from $\alpha$ by fixing $c$ as the minimum one with this property. Recall that $\mathfrak{q}_{1}^{q}\mathcal{O}_{L}\cdots\mathfrak{q}_{m}^{q}\mathcal{O}_{L}$ divides the ideal $(n+\alpha)\mathfrak{a}\mathcal{O}_{L}$ by (5.3) due to $\mathfrak{p}^{v}\mid\mathfrak{b}_{n}$. Then we obtain
$\mathfrak{q}_{1}^{q}\mathcal{O}_{L}\cdots\mathfrak{q}_{m}^{q}\mathcal{O}_{L}\mid(nc+\beta)\mathcal{O}_{L}$
since $(n+\alpha)\mathfrak{a}$ divides $(nc+\beta)$. In other words, $nc+\beta$ belongs to $\mathfrak{q}_{j}^{q}\mathcal{O}_{L}$ for any $j$. Suppose that $\mathcal{O}_{L}=\mathcal{O}_{M_{j}}[x_{1},\ldots,x_{d}]$ with $x_{1},\ldots,x_{d}\in\mathcal{O}_{L}$. Then we have
$1=a_{1}x_{1}+\cdots+a_{d}x_{d}\quad\text{and}\quad\beta=b_{1}x_{1}+\cdots+b_{d}x_{d}$
with some constants $a_{\ell},b_{\ell}\in\mathcal{O}_{M_{j}}$. It yields the formula
$nc+\beta=(na_{1}c+b_{1})x_{1}+\cdots+(na_{d}c+b_{d})x_{d}.$
Then $nc+\beta\in\mathfrak{q}_{j}^{q}\mathcal{O}_{L}$ is equivalent that
$(na_{1}c+b_{1})x_{1}+\cdots+(na_{d}c+b_{d})x_{d}=\sum_{t=1}^{T}p_{t}z_{t}$
with some $p_{t}\in\mathfrak{q}_{j}^{q}$ and $z_{t}\in\mathcal{O}_{L}$. We have
$\sum_{t=1}^{T}p_{t}z_{t}=\sum_{t=1}^{T}p_{t}(A_{t1}x_{1}+\cdots+A_{td}x_{d}),$
where $A_{t1},\ldots,A_{td}\in\mathcal{O}_{M_{j}}$. Comparing them, we derive
$na_{\ell}c+b_{\ell}=\sum_{t=1}^{T}p_{t}A_{t\ell}\in\mathfrak{q}_{j}^{q}$
for any $\ell\in\{1,\ldots,d\}$. This implies that $a_{k}b_{\ell}-a_{\ell}b_{k}\in\mathfrak{q}_{j}^{q}$ for any $k,\ell\in\{1,\ldots,d\}$. Hence we conclude that $\mathfrak{q}_{j}^{q}$ divides the integral ideal generated by all elements in $\{a_{k}b_{\ell}-a_{\ell}b_{k}\mid k\neq\ell\}$ over $\mathcal{O}_{M_{j}}$. Define $G(\alpha;M_{j})$ as the norm of this ideal. One can check that $G(\alpha;M_{j})$ is independent to the choice of the basis $\{x_{1},\ldots,x_{d}\}$. Furthermore, $G(\alpha)=\max G(\alpha;M)$ depends only on $\alpha$, where the maximum is taken over all algebraic fields $M$ such that $\mathbb{Q}\subset M\subsetneq L$. Finally, we see that
(5.6) $\operatorname{N}(\mathfrak{q}_{1}^{q})^{d}\cdots\operatorname{N}(\mathfrak{q}_{m}^{q})^{d}\leq G(\alpha)^{dm}\leq G(\alpha)^{[L:\mathbb{Q}]}$
due to $d=ef$ and $[L:\mathbb{Q}]=efg$. By (5.4), (5.5), and (5.6), we arrive at
$\operatorname{N}(\mathfrak{p}^{v})\leq\left(G(\alpha)\Delta_{L/\mathbb{Q}}\right)^{[K:\mathbb{Q}]}=C_{1}(\alpha),$
say. Then the desired result immediately follows. Indeed, we have
$\operatorname{N}(\mathfrak{b}_{n})=\prod_{\mathfrak{p}\mid\mathfrak{b}_{n}}\operatorname{N}(\mathfrak{p}^{\nu_{\mathfrak{p}}(\mathfrak{b}_{n})})\leq C_{1}(\alpha)^{\pi_{K}(C_{1}(\alpha))},$
where $\pi_{K}(x)$ denotes the number of prime ideals $\mathfrak{p}$ of $K$ such that $\operatorname{N}(\mathfrak{p})\leq x$. ∎
Lemma 5.2.
@@
Take an irreducible polynomial $f(t)\in\mathbb{Z}[t]$ with $f(-\alpha)=0$, and denote by $A_{f}$ its leading coefficient. Let $C(\alpha)>0$ be the constant described in Lemma 5.1. Then, for any $n\in\mathbb{Z}$ and any $x>\max\{C(\alpha),\operatorname{N}(\mathfrak{a}),|A_{f}|\}$, the following conditions are equivalent.
@@
    There exists a prime ideal $\mathfrak{p}\in P_{K}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})>x$.
@@
holds, where $\alpha_{1},\ldots,\alpha_{d}$ are the conjugates of $\alpha$ over $\mathbb{Q}$. Suppose that there exists a rational prime $p$ such that $p\mid f(n)$ and $p>x$. By (5.7), we deduce that $p$ divides $\operatorname{N}((n+\alpha)\mathfrak{a})$ or $A_{f}$, but the later does not hold due to $x>|A_{f}|$. Thus we have that $p\mid\operatorname{N}((n+\alpha)\mathfrak{a})$. Furthermore, it implies that there exists a prime ideal $\mathfrak{p}$ of $K$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})=p^{f}$ with some $j\geq 1$. If we suppose that $\mathfrak{p}\notin P_{K}$, then $\operatorname{N}(\mathfrak{p})>x>C(\alpha)$ would follow, which yields contradiction by Lemma 5.1. Hence we conclude that $\mathfrak{p}\in P_{K}$ holds, and therefore, $(\mathrm{A})$ implies $(\mathrm{B})$.
Suppose conversely that there exists a prime ideal $\mathfrak{p}\in P_{K}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})>x$. Then $\operatorname{N}(\mathfrak{p})=p$ holds for a rational prime $p$ by the definition of $P_{K}$. Thus $p\mid\operatorname{N}((n+\alpha)\mathfrak{a})$ and $p>x$ are satisfied. Furthermore, $p\mid\operatorname{N}((n+\alpha)\mathfrak{a})$ implies that $p$ divides $f(n)$ or $\operatorname{N}(\mathfrak{a})$ by (5.7), but the later does not hold due to $x>\operatorname{N}(\mathfrak{a})$. As a result, we conclude that $(\mathrm{B})$ implies $(\mathrm{A})$. ∎
Lemma 5.3.
Let $\mathfrak{p}\in P_{K}$ and $v\geq 1$. Assume $\mathfrak{p}^{v}\mid(m+\alpha)\mathfrak{a}$ and $\mathfrak{p}^{v}\mid(n+\alpha)\mathfrak{a}$ for $m,n\in\mathbb{Z}$. Then we have $m\equiv n~{}(\bmod\,{p^{v}})$ with $\operatorname{N}(\mathfrak{p})=p$.
Proof.
Put $m-n=p^{w}A$ with some $w\in\mathbb{Z}_{\geq 0}$ and $A\in\mathbb{Z}_{\geq 1}$ such that $(p,A)=1$. By the assumption, we have $(m+\alpha)\mathfrak{a}\subset\mathfrak{p}^{v}$ and $(n+\alpha)\mathfrak{a}\subset\mathfrak{p}^{v}$. Thus we see that $(m-n)\mathfrak{a}\subset\mathfrak{p}^{v}$ holds, which yields
$\mathfrak{p}^{v}\mid(m-n)$
due to $\mathfrak{p}\nmid\mathfrak{a}$. Therefore, $\mathfrak{p}^{v}\mid(p)^{w}(A)$ follows. We deduce that $\mathfrak{p}\nmid(A)$ from $(p,A)=1$ by taking the norms. As a result, we obtain
$\mathfrak{p}^{v}\mid(p)^{w}.$
Recall that $\mathfrak{p}$ is unambiguous, i.e. $\mathfrak{p}^{2}\nmid(p)$. Hence we see that $w\geq v$, which implies $p^{v}\mid(m-n)$ as desired. ∎
From the above, we associate $C_{\alpha}(x)$ defined in Section 1.2 with $\Psi_{f}(x,y)$ for an irreducible polynomial $f(t)\in\mathbb{Z}[t]$ with $f(-\alpha)=0$. The relation between $C_{\alpha}(x)$ and $\Psi_{f}(x,y)$ was mentioned firstly by Worley [20], and a more detailed research was presented by Andersson in his unpublished work [1]. In this paper, we obtain the following result.
Proposition 5.4.
$C_{\alpha}(x)=x-\Psi_{f}(x,x)+O\left(\frac{x}{\log{x}}\right)$
Proof.
@@
with an absolute implied constant. Take an integer $n\in N_{f}(x)$ arbitrarily. Then there exists a rational prime $p$ such that $p\mid f(n)$ and $p>x$. Here, we suppose $x>\max\{C(\alpha),\operatorname{N}(\mathfrak{a}),|A_{f}|\}$ as in Lemma 5.2. Then we deduce from Lemma 5.2 that there exists a prime ideal $\mathfrak{p}\in P_{K}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})>x$. If $\mathfrak{p}\mid(m+\alpha)\mathfrak{a}$ is satisfied for some $m\in\mathbb{Z}$ in $[1,x]$, then we apply Lemma 5.3 to deduce that $m\equiv n~{}(\bmod\,{p})$ with $\operatorname{N}(\mathfrak{p})=p$. Since $m,n\in[1,x]$ and $p>x$, we must obtain $m=n$, and therefore, the condition
@@
Proof of Theorem 1.2.
Let $f(t)\in\mathbb{Z}[t]$ be an irreducible polynomial with $f(-\alpha)=0$. Then $f(t)$ is of degree $d=\deg(\alpha)\geq 2$. By Proposition 5.4 and Theorem 1.1 with $u=1$, we have
@@
$0.543\displaystyle 1-\frac{\gamma_{f}(1)}{d}>1-\frac{0.914}{2}=$ 0.543
@@
Let $f(t)$ be an irreducible polynomial with $f(-\alpha)=0$. If (1.1) is true at least for $u=1$, then we deduce from Proposition 5.4 that
@@
Remark 5.5.
Actually, Lemma 5.1 is much stronger than necessary for the proof of Theorem 1.2 since one needs just that $\operatorname{N}(\mathfrak{p})$ is bounded for any $\mathfrak{p}\notin P_{K}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$. We note that Lemma 5.1 is a modification of a statement of Cassels [2]. He used the set $P_{K}^{*}$ consisting of prime ideals $\mathfrak{p}$ of $K$ such that
  • $(\mathrm{i})$
    $\mathfrak{p}$ is of the first degree;
  • $(\mathrm{ii})$
    $\mathfrak{p}$ is unambiguous;
  • $(\mathrm{iii})$
    for any $m\in\mathbb{Z}$, if $\mathfrak{p}\mid(m+\alpha)\mathfrak{a}$ then $(m+\alpha)\mathfrak{a}$ is not divisible by any prime ideal $\mathfrak{p}^{\prime}\neq\mathfrak{p}$ satisfying $\operatorname{N}(\mathfrak{p})=\operatorname{N}(\mathfrak{p}^{\prime})$.
We factorize the ideal $(n+\alpha)\mathfrak{a}$ as $(n+\alpha)\mathfrak{a}=\mathfrak{a}_{n}^{*}\mathfrak{b}_{n}^{*}$, where
$\mathfrak{a}_{n}^{*}=\prod_{\mathfrak{p}\in P_{K}^{*}}\mathfrak{p}^{v_{\mathfrak{p}}(n+\alpha)}\quad\text{and}\quad\mathfrak{b}_{n}^{*}=\prod_{\mathfrak{p}\notin P_{K}^{*}}\mathfrak{p}^{v_{\mathfrak{p}}(n+\alpha)}.$
Then it was claimed in [2, p. 181] that $\operatorname{N}(\mathfrak{b}_{n}^{*})$ is bounded for $n\in\mathbb{Z}$. However, we know that $\operatorname{N}(\mathfrak{b}_{n}^{*})$ is not bounded in general. For example, letting $\alpha=7\sqrt{2}$, one can find for any $v\geq 1$ an integer $n=n(v)$ such that $\operatorname{N}(b_{n}^{*})\geq 7^{v}$. See [16, Remark 1.6] for more information. Hence the original method by Cassels seems flawed.
The boundedness of $\operatorname{N}(\mathfrak{b}_{n}^{*})$ was used in [2] to deduce the inequality
$\sum_{\mathfrak{p}\in P_{K}^{*}}v_{\mathfrak{p}}(n+\alpha)\log{\operatorname{N}(\mathfrak{p})}>2\log{n}-C$
for a constant $C>0$ which is independent of $n$. Furthermore, condition $(\mathrm{iii})$ was used to see that there is no contribution of terms for $\mathfrak{p}_{1},\mathfrak{p}_{2}$ such that $\mathfrak{p}_{1}\neq\mathfrak{p}_{2}$ but $\operatorname{N}(\mathfrak{p}_{1})=\operatorname{N}(\mathfrak{p}_{2})$ in the sum
(5.9) $\sum_{n\in\mathfrak{S}_{N}}\sum_{\begin{subarray}{c}\mathfrak{p}_{1},\mathfrak{p}_{2}\in P_{K}^{*}\\ \mathfrak{p}_{j}\mid(n+\alpha)\mathfrak{a}\\ \operatorname{N}(\mathfrak{p}_{j})<\sqrt{M}\end{subarray}}\frac{\log{\operatorname{N}(\mathfrak{p}_{1})}}{\operatorname{N}(\mathfrak{p}_{1})}\frac{\log{\operatorname{N}(\mathfrak{p}_{2})}}{\operatorname{N}(\mathfrak{p}_{2})},$
where $\mathfrak{S}_{N}$ is the set of integers $n$ in $[N,N+M]$ such that $\operatorname{N}(\mathfrak{p}^{v_{\mathfrak{p}}(n+\alpha)})<M$ holds for all $\mathfrak{p}\in P_{K}^{*}$ with $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$. Note that (5.9) appeared during a calculation using the Cauchy–Schwarz inequality, and that a similar argument was also presented in the proof of Proposition 2.1. The difference is that the method of this paper is non ideal-theoretic, which requires nothing similar to condition $(\mathrm{iii})$. On the other hand, it is also possible to recover the proof in [2] by replacing $P_{K}^{*}$ with $P_{K}$. Indeed, $\operatorname{N}(\mathfrak{b}_{n})$ is bounded by Lemma 5.1, which yields
$\sum_{\mathfrak{p}\in P_{K}}v_{\mathfrak{p}}(n+\alpha)\log{\operatorname{N}(\mathfrak{p})}>2\log{n}-C.$
Furthermore, for any prime ideals $\mathfrak{p}_{1},\mathfrak{p}_{2}\in P_{K}$ with $\mathfrak{p}_{j}\mid(n+\alpha)\mathfrak{a}$ such that $\mathfrak{p}_{1}\neq\mathfrak{p}_{2}$ but $\operatorname{N}(\mathfrak{p}_{1})=\operatorname{N}(\mathfrak{p}_{2})=:p$, one can show that $p$ is bounded for $n\in\mathbb{Z}$. Actually, this is sufficient to show that the contribution of terms for $\mathfrak{p}_{1},\mathfrak{p}_{2}$ such that $\mathfrak{p}_{1}\neq\mathfrak{p}_{2}$ but $\operatorname{N}(\mathfrak{p}_{1})=\operatorname{N}(\mathfrak{p}_{2})$ in (5.9) is not zero but can be negligible.
To close this section, we confirm the original statement of Cassels [2].
@@
Let $\alpha$ be an algebraic irrational number. Then there exists an integer $N_{0}=N_{0}(\alpha)>10^{6}$ depending only on $\alpha$ with the following property. Suppose that $N\geq N_{0}$ and put $M=\lfloor 10^{-6}N\rfloor$. Then at least $10051M/$ 100 integers in $N<n\leq N+M$ are such that $(n+\alpha)\mathfrak{a}$ is divisible by a prime ideal $\mathfrak{p}$ which does not divide $(m+\alpha)\mathfrak{a}$ for any integer $0\leq m\leq N+M$ with $m\neq n$.
@@
by (2.2) if $N$ is large enough. Furthermore, along the same line as the proof of Proposition 5.4, we see that the number of integers in $N<n\leq N+M$ with the desired property is equal to
@@
6. Proof of Theorems 1.3 and 1.4
We prove Theorems 1.3 and 1.4 by associating the functions $R_{b}(x)$ and $N(x)$ defined in Section 1.3 with $\Psi_{f}(x,y)$ for a suitable polynomial $f(t)$. For this, we recall the following criterions.
@@
Let $n\in\mathbb{Z}_{>|b|}$ with $-b$ not an integer square. Then $n^{2}+b$ has a primitive divisor if and only if $P^{+}(n^{2}+b)>2n$.
@@
Then the equality $N(x)=R_{1}(x)$ immediately follows. Furthermore, we associate $R_{b}(x)$ with $\Psi_{f}(x,y)$ for the quadratic polynomial $f(t)=t^{2}+b$ as follows.
Proposition 6.3.
Let $f(t)=t^{2}+b$ with $-b$ not an integer square. Then we have
$R_{b}(x)=x-\Psi_{f}(x,x)+O\left(\frac{x\log\log{x}}{\log{x}}\right)$
Proof.
@@
with a parameter $z$ such that $21\leq z\leq x/$ 2. From the above, we obtain the formula
@@
$\sum_{\begin{subarray}{c}z\leq n\leq x\\ 2z\leq P^{+}(n^{2}+b)<2x\end{subarray}}1=\sum_{2z\leq p<2x}\sum_{\begin{subarray}{c}z\leq n\leq x\\ P^{+}(n^{2}+b)=p\end{subarray}}1\leq\sum_{2z\leq p<2x}\sum_{\begin{subarray}{c}z\leq n\leq x\\ f(n)\equiv 0~{}\pmod{p}\end{subarray}}1$
@@
$\sum_{2z\leq p<2x}\sum_{\begin{subarray}{c}z\leq n\leq x\\ f(n)\equiv 0\pmod{p}\end{subarray}}1$ $\ll\sum_{2z\leq p<2x}\left(\frac{x}{p}+1\right)$
@@
If (1.1) is true for $f(t)=t^{2}+b$ and $u=1$, then Proposition 6.3 yields that
$R_{b}(x)\sim(1-\rho(2))x$
Proof of Theorems 1.3 and 1.4.
The proof is similar to that of Theorem 1.2. Let $f(t)=t^{2}+b$. By Proposition 6.3 and Theorem 1.1 with $u=1$, we have
@@
$0.543\displaystyle 1-\frac{\gamma_{f}(1)}{2}>1-\frac{0.914}{2}=$ 0.543

### v2

@@ 3. An inductive process
Take an irreducible polynomial $f(t)\in\mathbb{Z}[t]$ with $f(-\alpha)=0$, and denote by $A_{f}$ its leading coefficient. Let $C(\alpha)>0$ be the constant described in (5.1). Then, for any $n\in\mathbb{Z}$ and any $x>\max\{C(\alpha),\operatorname{N}(\mathfrak{a}),|A_{f}|\}$, the following conditions are equivalent.
@@
    There exists a prime ideal $\mathfrak{p}\in P_{\alpha}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})>x$.
@@
holds, where $\alpha_{1},\ldots,\alpha_{d}$ are the conjugates of $\alpha$ over $\mathbb{Q}$. Suppose that there exists a rational prime $p$ such that $p\mid f(n)$ and $p>x$. By (5.2), we deduce that $p$ divides $\operatorname{N}((n+\alpha)\mathfrak{a})$ or $A_{f}$, but the later does not hold due to $x>|A_{f}|$. Thus we have that $p\mid\operatorname{N}((n+\alpha)\mathfrak{a})$. Furthermore, it implies that there exists a prime ideal $\mathfrak{p}$ of $K$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})=p^{f}$ with some $f\geq 1$. If we suppose that $\mathfrak{p}\notin P_{\alpha}$, then $\operatorname{N}(\mathfrak{p})>x>C(\alpha)$ would follow, which yields contradiction. Hence we conclude that $\mathfrak{p}\in P_{\alpha}$ holds, and therefore, $(\mathrm{A})$ implies $(\mathrm{B})$.
Suppose conversely that there exists a prime ideal $\mathfrak{p}\in P_{\alpha}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})>x$. Then $\operatorname{N}(\mathfrak{p})=p$ holds for a rational prime $p$ by the definition of $P_{\alpha}$. Thus $p\mid\operatorname{N}((n+\alpha)\mathfrak{a})$ and $p>x$ are satisfied. Furthermore, $p\mid\operatorname{N}((n+\alpha)\mathfrak{a})$ implies that $p$ divides $f(n)$ or $\operatorname{N}(\mathfrak{a})$ by (5.2), but the later does not hold due to $x>\operatorname{N}(\mathfrak{a})$. As a result, we conclude that $(\mathrm{B})$ implies $(\mathrm{A})$. ∎
Lemma 5.2 (Lemma 4.6 in [16]).
Let $\mathfrak{p}\in P_{\alpha}$ and $v\geq 1$. Assume $\mathfrak{p}^{v}\mid(m+\alpha)\mathfrak{a}$ and $\mathfrak{p}^{v}\mid(n+\alpha)\mathfrak{a}$ for $m,n\in\mathbb{Z}$. Then we have $m\equiv n~(\bmod\,{p^{v}})$ with $\operatorname{N}(\mathfrak{p})=p$.
Proof of Theorem 1.2.
@@
with an absolute implied constant. Take an integer $n\in N_{f}(x)$ arbitrarily. Then there exists a rational prime $p$ such that $p\mid f(n)$ and $p>x$. Here, we suppose $x>\max\{C(\alpha),\operatorname{N}(\mathfrak{a}),|A_{f}|\}$ as in Lemma 5.1. Then we deduce from Lemma 5.1 that there exists a prime ideal $\mathfrak{p}\in P_{\alpha}$ such that $\mathfrak{p}\mid(n+\alpha)\mathfrak{a}$ and $\operatorname{N}(\mathfrak{p})>x$. If $\mathfrak{p}\mid(m+\alpha)\mathfrak{a}$ is satisfied for some $m\in\mathbb{Z}$ in $[1,x]$, then we apply Lemma 5.2 to deduce that $m\equiv n~(\bmod\,{p})$ with $\operatorname{N}(\mathfrak{p})=p$. Since $m,n\in[1,x]$ and $p>x$, we must obtain $m=n$, and therefore, the condition
@@
Proof of Corollary 1.3.
Let $f(t)\in\mathbb{Z}[t]$ be an irreducible polynomial with $f(-\alpha)=0$. Then $f(t)$ is of degree $d=\deg(\alpha)\geq 2$. By Theorem 1.1 with $u=1$, we deduce from Theorem 1.2 that
@@
$1-\frac{\gamma_{f}(1)}{d}>1-\frac{0.914}{2}=0.543$
@@
Let $f(t)$ be an irreducible polynomial with $f(-\alpha)=1$. If (1.1) is true at least for $u=1$, then we deduce from Theorem 1.2 that
@@
Remark that the formulation of Corollary 1.3 is due to Worley [21], and the original statement by Cassels in [2] is slightly different. To close this section, we confirm the original statement of Cassels.
@@
Let $\alpha$ be an algebraic irrational number. Then there exists an integer $N_{0}=N_{0}(\alpha)>10^{6}$ depending only on $\alpha$ with the following property. Suppose that $N\geq N_{0}$ and put $M=\lfloor 10^{-6}N\rfloor$. Then at least $51M/100$ integers in $N<n\leq N+M$ are such that $(n+\alpha)\mathfrak{a}$ is divisible by a prime ideal $\mathfrak{p}$ which does not divide $(m+\alpha)\mathfrak{a}$ for any integer $0\leq m\leq N+M$ with $m\neq n$.
@@
by (2.2) if $N$ is large enough. Furthermore, along the same line as the proof of Theorem 1.2, we see that the number of integers in $N<n\leq N+M$ with the desired property is equal to
@@
6. Proof of Theorem 1.4 and its corollaries
Firstly, we recall the following criterions, which yield $N(x)=R_{1}(x)$ immediately.
@@
Let $n\in\mathbb{Z}_{>|b|}$ with b-b not an integer square. Then $n^{2}+b$ has a primitive divisor if and only if $P^{+}(n^{2}+b)>2n$.
@@
Furthermore, Theorem 1.4 is proved by using Lemma 6.1 as follows.
Proof of Theorem 1.4.
@@
with a parameter $z$ such that $1\leq z\leq x/2$. From the above, we obtain the formula
@@
$\sum_{\begin{subarray}{c}z\leq n\leq x\\ 2z\leq P^{+}(n^{2}+b)<2x\end{subarray}}1=\sum_{2z\leq p<2x}\sum_{\begin{subarray}{c}z\leq n\leq x\\ P^{+}(n^{2}+b)=p\end{subarray}}1\leq\sum_{2z\leq p<2x}\sum_{\begin{subarray}{c}z\leq n\leq x\\ f(n)\equiv 0~\hskip-5.69054pt\pmod{p}\end{subarray}}1$
@@
$\sum_{2z\leq p<2x}\sum_{\begin{subarray}{c}z\leq n\leq x\\ f(n)\equiv 0\hskip-5.69054pt\pmod{p}\end{subarray}}1$ $\ll\sum_{2z\leq p<2x}\left(\frac{x}{p}+1\right)$
@@
Proof of Corollaries 1.5 and 1.6.
The proof is similar to that of Corollary 1.3. Let $f(t)=t^{2}+b$. By Theorem 1.1 with $u=1$, we deduce from Theorem 1.4 that
@@
$1-\frac{\gamma_{f}(1)}{2}>1-\frac{0.914}{2}=0.543$